OpenStudy (anonymous):

Find k such that y=k-x^2 is tangent to y=-6x+7

OpenStudy (anonymous):

solve both the equations

OpenStudy (anonymous):

i know the derivative of the second one is -6 but i dont know where to go from there

OpenStudy (anonymous):

Since the slope of the tangent line is -6, set the derivative of the first equation to -6... ie, -2x = -6 x = ???

OpenStudy (anonymous):

^^ that is the x-coordinate where tangency occurs.

OpenStudy (anonymous):

x=3 is not the answer though. i just tried it. they are looking for k

OpenStudy (anonymous):

k=-2 right? (trying to follow along)

OpenStudy (anonymous):

I know... we're not done yet....

OpenStudy (anonymous):

yes it is -2! how did you find that

OpenStudy (anonymous):

At that x coordinate, find the y-coordinate of the point of tangency.... so, y = -6(3) + 7 y = ????

OpenStudy (anonymous):

(and that will not be the answer yet either...)

OpenStudy (anonymous):

should give them the solution as you go so they know what they should be getting y=k - x^2

OpenStudy (anonymous):

Did you find y = ???

OpenStudy (anonymous):

y =-11

OpenStudy (anonymous):

Ok..., so plugging this into the first equation along with x=3, you have: -11 = k -(3)^2 now do you know how k is obtained?

OpenStudy (anonymous):

ohh okay yes thank you!

OpenStudy (anonymous):

you're welcome